# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

### v1

@@ Abstract
In this paper, we propose a Sparse Low-rank Quaternion Approximation (SLRQA) model for color image processing problems with noisy observations. The proposed SLRQA is a quaternion model that combines low-rankness and sparsity priors. Furthermore, it does not need an initial rank estimate. To solve the SLRQA model, a nonconvex proximal linearized ADMM (PL-ADMM) algorithm is proposed. Furthermore, the global convergence analysis of the PL-ADMM under mild assumptions is presented. When the observation is noise-free, an SLRQA-NF model of the limiting case of the SLRQA is proposed. Subsequently, a nonconvex proximal linearized ADMM (PL-ADMM-NF) algorithm for the SLRQA-NF is given. In numerical experiments, we verify the effectiveness of quaternion representation. Furthermore, for color image denoising and color image inpainting problems, SLRQA and SLRQA-NF models demonstrate superior performance both quantitatively and visually when compared with some state-of-the-art methods.
@@ 1 Introduction
Low-rankness plays an important role in image processing and has been used in image denoising [10], image inpainting [15], image deblurring [26], and image filtering [7]. A commonly encountered approach for exploiting the low-rankness is to find a matrix with the lowest rank such that certain matching errors are minimized. Due to the existence of the rank function, such problems are known as NP-hard [25]. To overcome this difficulty, Candès et al. [6] have proved that the nuclear norm is the tightest convex relaxation of the NP-hard rank minimization function. In addition to the nuclear norm, other surrogates have also been proposed, such as weighted nuclear norm (WNNM) [14], Schatten $p$-norm [24], weighted Schatten $p$-norm [32], and log-determinant penalty [19], which have shown competitive effectiveness in various applications. In addition to low-rankness, sparsity is also a significant property of natural images that has been considered in image recovery [35, 5], video recovery [34], texture repairing [21], and face recognition [40]. The intuition behind sparsity is that images under certain transforms are usually sparse, such as wavelet or discrete cosine transform (DCT). Moreover, quaternion representation has shown great potential for color image processing problems. A color image can be encoded as a pure quaternion matrix. As claimed in [10], the main advantage of quaternion representation is that a color image is treated as a pure quaternion matrix naturally and holistically, meaning that the correlation information among the three RGB channels is used. It is empirically shown in [10] that methods using quaternion representation outperform methods with monochromatic techniques. Moreover, quaternion has been increasingly used for color image processing in the past two decades [10].
These three priors, i.e. low-rankness, sparsity, and quaternion representation, all play important roles in image processing. However, models based on all these three priors are still limited. In the following subsection, we propose a quaternion-based model that combines low-rankness and sparsity properties to address problems arising in image processing.
@@ 1.1 A novel quaternion based model
By integrating the low-rankness and sparsity into the prior term, i.e., $R({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathcal{W}({\mathbf{\dot{X}}}))$, and letting $\mathbf{\dot{W}}=\mathcal{W}({\mathbf{\dot{X}}})$, we propose the following Sparse Low-rank Quaternion Approximation (SLRQA) model for color image processing:
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2},$ (2)
@@
where ${\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}$ denotes a color image, $\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}$ is the coefficient matrix under an orthogonal transform $\mathcal{W}$ such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT), $\mathcal{A}$ is a known operator such as a total variation operator, a blurring operator or the Radon transform, $p(\mathbf{\dot{W}})$ is a function that has Lipschitz continuously differentiable and promotes the sparsity of $\mathbf{\dot{W}}$, $\phi$ is a nonconvex surrogate function that substitutes the rank function, $\sigma_{i}({\mathbf{\dot{X}}})$ is the $i$-th singular value of ${\mathbf{\dot{X}}}$ defined in Section 2, $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ promotes the low-rankness of ${\mathbf{\dot{X}}}$. Some nonconvex surrogate functions for $\phi$ have been given in [10, 23] and listed in Table 1.
Table 1: nonconvex surrogate functions $\phi(x,\gamma)$
@@
The $\ell_{1}$ norm has been widely used to promote sparsity. However, the Lipschitz continuous differentiability of $p(\mathbf{\dot{W}})$ is assumed in our convergence analysis [^2] We use a variant of the ADMM method to solve SLRQA. Note that such an assumption has been made for all the ADMM methods with global convergence guaranteed for nonconvex problems as far as we know, see, e.g., a review in [22]. Therefore, a smoothing function such as the Huber function can be used as a substitution for $p(\mathbf{\dot{W}})$. It is empirically shown in Section 6 that such a substitution yields satisfactory performance. This strategy has been used in [38] to guarantee convergence theoretically.
@@
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (3)
@@
Since $\mathcal{W}$ is an orthogonal transform such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT) [12], we have $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Q}}_{1}\mathbf{\dot{W}}\mathbf{\dot{Q}}_{2}^{*}),\gamma)=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{W}}),\gamma)$, the superscript * denotes the conjugate transpose operator. Let $\mathbf{\dot{Z}}=\mathcal{W}({\mathbf{\dot{X}}})$. We propose a model named SLRQA-NF which is equivalent to (3) for noise-free problems:
$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (4)
@@ 1.2 Related work
The derivations of SLRQA and SLRQA-NF models are motivated by LADM [21] and LRQA [10] models. To be specific, LADM is a convex model that utilizes the sparsity and low-rankness properties while quaternion representation and nonconvex surrogate rank function have not been considered. LRQA is a quaternion-based model that makes use of the low-rankness property by replacing the nuclear norm with nonconvex surrogate functions, but it has not taken advantage of sparsity. SLRQA and SLRQA-NF models take advantage of LADM and LRQA by using quaternion representation, low-rankness, and sparsity. Next, we discuss some image processing models related to SLRQA and SLRQA-NF.
@@
where the variable $\mathbf{\dot{W}}$ is eliminated by $\mathbf{\dot{W}}=\mathcal{W}({\mathbf{\dot{X}}})$. If the sparse term $p(\mathbf{\dot{W}})$ is not used, then SLRQA becomes the LRQA denoising model in [10, Equation 20]
@@
As noted in [10], LRQA is a generalization of other models such as the WNNM model.
@@
Image inpainting (partially observed images): If the observed image is noise-free, the linear operator $\mathcal{A}$ is the identity, $p(\mathbf{\dot{W}})=\|\mathbf{\dot{W}}\|_{1}$, and $\phi$ is the nuclear norm, then the SLRQA-NF model becomes the LADM model in [21] with an extension to quaternion representation:
@@
Furthermore, if the sparsity term $\phi$ is dropped and a nonconvex surrogate is used, then SLRQA-NF model becomes the LRQA [10, Equation 22] given by
@@
where $r$ is a prescribed number and the superscript * denotes the conjugate transpose operator, see its definition in Section 2. Different from SLRQA models, Model (8) uses the nuclear norm to promote low-rank instead of nonconvex surrogate functions $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$. Furthermore, it needs to estimate the rank $r$ in the initialization phase while SLRQA does not. Another model that uses the three priors is proposed in [34, 35] for color image and video recovery.
The model is given by
$\min_{{\mathbf{\dot{X}}}}$ $\|{\mathbf{\dot{X}}}\|_{*}-\sum_{i=r}^{\min(m,n)}\sigma_{i}({\mathbf{\dot{X}}})+\lambda\|\mathbf{\dot{W}}\|_{1},$ (9)
@@
Note that Model (9) also needs to estimate the rank $r$ and the low-rankness is promoted by penalizing a truncated nuclear norm, which is convex. Moreover, convergence analyses of the proposed algorithms for solving the existing models in [17, 34, 35] with three priors are not given.
In this paper, proximal linearized ADMM algorithms are proposed to solve Models (2) and (4). Note that the convergence of nonconvex ADMM and its variants has been studied widely. However, as far as we know, the existing ADMM and its variants either consider different optimization problems or use different optimization techniques. A nonexhaustive review is given as follows. Wang et al. [29] comprehensively study the convergence of ADMM for nonconvex and nonsmooth objective functions subject to coupled linear equality constraints. Their assumptions require that the solution of the subproblem is Lipschitz continuous with respect to the input, which is often overly stringent and hard to verify for practical applications. Hong et al. [18] study the convergence of ADMM for solving nonconvex consensus problems, but it is under the case where every constraint matrix has full column rank. Compared with the above analyses, our convergence analysis only needs one constraint matrix that has full column rank. In [36], Yashtini proposes a proximal linearized ADMM for solving linearly constrained nonconvex and possibly nonsmooth optimization problems. The proposed method employs a linearization strategy that targets both the quadratic penalty term and the objective function. Above all, all the nonconvex ADMM and their variant algorithms for nonconvex situations are different from PL-ADMM where only one variable is linearized in the situation. For the reviewed algorithms, the first two algorithms are traditional ADMM algorithms for nonconvex problems, while the ADMM of the last algorithm uses the technique of linearization for both two variables. For a detailed review of other non-convex ADMM algorithms, we refer the readers to [22].

### v2

@@ Abstract
In this paper, we propose a Sparse Low-rank Quaternion Approximation (SLRQA) model for color image processing problems with noisy observations. The proposed SLRQA is a quaternion model that combines low-rankness and sparsity priors without an initial rank estimation. A proximal linearized ADMM (PL-ADMM) algorithm is proposed to solve SLRQA and the global convergence is guaranteed under standard assumptions. When the observation is noise-free, a limiting case of the SLRQA, called SLRQA-NF, is proposed. Subsequently, a proximal linearized ADMM (PL-ADMM-NF) algorithm for SLRQA-NF is given. Since SLRQA-NF does not satisfy a widely-used assumption for global convergence of ADMM-type algorithms, we propose a novel assumption, under which the global convergence of PL-ADMM-NF is established. In numerical experiments, we verify the effectiveness of quaternion representation. Furthermore, for color image denoising and color image inpainting problems, SLRQA and SLRQA-NF demonstrate superior performance both quantitatively and visually when compared with some state-of-the-art methods.
@@ 1 Introduction
Low-rankness plays an important role in image processing and has been used in image denoising [12], image inpainting [16], image deblurring [31], and image filtering problems [9]. A commonly encountered approach for exploiting the low-rankness is to find a matrix with the lowest rank such that certain matching errors are minimized. Due to the existence of the rank function, such problems are known as NP-hard [30]. To overcome this difficulty, Candès et al. have proved that the nuclear norm is the tightest convex relaxation of the NP-hard rank minimization function [8]. In addition to the nuclear norm, other surrogates have also been proposed, such as weighted nuclear norm (WNNM)[15], Schatten $p$-norm [29], weighted Schatten $p$-norm [38], and log-determinant penalty [21], which have shown competitive effectiveness in various applications. In addition to low-rankness, sparsity is also a significant property of natural images that has been considered in image recovery [42, 7], video recovery [41], texture repairing [26], and face recognition [50]. The intuition behind sparsity is that images under certain transforms are usually sparse, such as wavelet or discrete cosine transform (DCT). Moreover, quaternion representation has shown great potential for color image processing problems. A color image can be encoded as a pure quaternion matrix. The main advantage of quaternion representation is that the three color channels simultaneously are treated by quaternion representation, which can well keep the connection between image color channels [39]. Moreover, compared to tensor-based methods, quaternion-based methods not only use the information among channels but also preserve the orthogonality of the coefficient matrix for the three channels, see detailed discussions under the framework of dictionary learning in [40, Section III]. Furthermore, it is empirically shown in [40, 12] that quaternion representation consistently provides a more accurate approximation than monochromatic representation using the reconstructed matrix of the same rank.
The three priors, i.e. low-rankness, sparsity, and quaternion representation, all play important roles in image processing. However, models based on all these three priors are still limited. In the following subsection, we propose a quaternion-based model that combines low-rankness and sparsity properties to address problems arising in image processing.
@@ 1.1 A novel quaternion based model
By integrating the low-rankness and sparsity into the prior term, i.e., $R({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathcal{W}({\mathbf{\dot{X}}}))$, and letting $\mathbf{\dot{W}}=\mathcal{W}({\mathbf{\dot{X}}})$, we propose the following Sparse Low-rank Quaternion Approximation (SLRQA) model for color image processing:
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2},$ (2)
@@
where ${\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}$ denotes a color image, $\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}$ is the coefficient matrix under an orthogonal transform $\mathcal{W}$ such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT), $\mathcal{A}$ is a known operator such as a blurring operator or the Radon transform, $p(\mathbf{\dot{W}})$ is a function that promotes sparsity and has Lipschitz continuously differentiable such as a Huber function, i.e., $(p(\mathbf{\dot{W}}))_{ij}=\frac{1}{2\delta}|\mathbf{\dot{W}}_{ij}|^{2}$ if $|\mathbf{\dot{W}}_{ij}|<\delta$; and $(p(\mathbf{\dot{W}}))_{ij}=|\mathbf{\dot{W}}_{ij}|-\frac{1}{2}\delta$ otherwise, $\delta>0$ is a prescribed constant, $\varepsilon>0$ is a smoothing parameter, $\phi$ is a nonconvex surrogate function that substitutes the rank function, $\sigma_{i}({\mathbf{\dot{X}}})$ is the $i$-th singular value of ${\mathbf{\dot{X}}}$ defined in Section 2. The function $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$ promotes the low-rankness of ${\mathbf{\dot{X}}}$. We note that if $\varepsilon=0$, then $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$ reduces to $\sum_{i}\phi(\sigma_{i}({{\mathbf{\dot{X}}}}),\gamma)$. If we futher assumes that $\phi(x,\gamma)=x$, then $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$ reduces to the nuclear norm. Some nonconvex surrogate functions for $\phi$ have been given in [12, 28] and listed in Table 1.
Table 1: Some surrogate functions $\phi(x,\gamma)$
@@
The $\ell_{1}$ norm and nuclear norm have been widely used to promote sparsity and low rankness respectively. However, the Lipschitz continuous differentiability of $p(\mathbf{\dot{W}})$ is assumed in our convergence analysis [^2] We use a variant of the ADMM method to solve SLRQA. Note that such an assumption has been made for all the ADMM methods with global convergence guaranteed for nonconvex problems as far as we know, see, e.g., a review in [27].. Therefore, a smoothing function such as the Huber function can be used as a substitution for $p(\mathbf{\dot{W}})$. For more surrogate function of $\|\cdot\|_{1}$, we refer the readers to [28]. For the subsequent convergence analysis for Model (4), it is required that the term that promotes the low-rankness is also differentiable with respect to ${\mathbf{\dot{X}}}$. To satisfy this condition and ensure numerical stability, we also use a smoothed version of $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ by adding a small positive perturbation $\varepsilon$, i.e., replacing $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ with $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma),$ as discussed in [46, 47]. It is empirically shown in Section 6 that such a substitution yields satisfactory performance. This strategy has also been used in [46, 47] to guarantee convergence theoretically.
@@
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (3)
@@
Since $\mathcal{W}$ is an orthogonal transform such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT) [14], we have $\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{X}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)=\sum_{i}\phi(\sigma_{i}(\sqrt{\mathbf{\dot{W}}^{*}\mathbf{\dot{W}}+\varepsilon^{2}\mathbf{I}}),\gamma)$, where $\mathbf{\dot{Q}}_{1},\mathbf{\dot{Q}}_{2}$ are unitary matrix that defined later, the superscript * denotes the conjugate transpose operator. Let $\mathbf{\dot{Z}}=\mathcal{W}({\mathbf{\dot{X}}})$. We propose a model named SLRQA-NF which is equivalent to (3) for noise-free problems:
$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (4)
@@ 1.2 Related work
The derivations of SLRQA and SLRQA-NF are motivated by sparse low rank inpainting (SLRI) [26] and Low-Rank Quaternion Approximation (LRQA) [12] models. To be specific, SLRI is a convex model that uses the sparsity and low-rankness properties while quaternion representation and nonconvex surrogate rank function have not been considered. LRQA is a quaternion-based model that makes use of the low-rankness property by replacing the nuclear norm with nonconvex surrogate functions, but it has not taken advantage of sparsity. SLRQA and SLRQA-NF take advantage of SLRI and LRQA by using quaternion representation, low-rankness, and sparsity. Next, we discuss some image processing models related to SLRQA and SLRQA-NF.
@@
where the variable $\mathbf{\dot{W}}$ is eliminated by $\mathbf{\dot{W}}=\mathcal{W}({\mathbf{\dot{X}}})$. If the sparse term $p(\mathbf{\dot{W}})$ is not used, the low-rankness prior is used and $\varepsilon$ is set to 0, then SLRQA becomes LRQA denoising model in [12, Equation (20)]
@@
As noted in [12], LRQA is a generalization of other models such as the WNNM [15].
@@
Image inpainting (partially observed images): If the observed image is noise-free, the linear operator $\mathcal{A}$ is the identity, $p(\mathbf{\dot{W}})=\|\mathbf{\dot{W}}\|_{1}$, $\phi$ is the nuclear norm and $\varepsilon=0$, then SLRQA-NF becomes SLRI in [26] with an extension to quaternion representation:
@@
Furthermore, if the sparsity term is dropped, a nonconvex surrogate $\phi$ is used and $\varepsilon$ is set to 0, then SLRQA-NF becomes LRQA [12, Equation (22)] given by
@@
where $r$ is a prescribed integer and the superscript * denotes the conjugate transpose operator, see its definition in Section 2. Different from SLRQAs, Model (8) uses the nuclear norm to promote low-rank instead of nonconvex surrogate functions $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$. Furthermore, it needs to estimate the rank $r$ in the initialization phase while SLRQA does not. Another model that uses the three priors is proposed in [41, 42] for color image and video recovery. The model is given by
$\min_{{\mathbf{\dot{X}}}\in\mathbb{R}^{m\times n}}$ $\|{\mathbf{\dot{X}}}\|_{*}-\sum_{i=r}^{\min(m,n)}\sigma_{i}({\mathbf{\dot{X}}})+\lambda\|\mathbf{\dot{W}}\|_{1},$ (9)
@@
Note that Model (9) also needs to estimate the rank $r$ and the low-rankness is promoted by penalizing a truncated nuclear norm, which is nonconvex. Moreover, convergence analysis of the proposed algorithms for solving the existing models in [18, 41, 42] with three priors are not given.
In this paper, proximal linearized ADMM algorithms, i.e., PL-ADMM and PL-ADMM-NF are proposed to solve Models (2) and (4) and the global convergence of these algorithms is established. A generalization of SLRQA is formulated as
where $f$ is semi-lowercontinuous, $g$ is continuously differentiable, $\mathcal{C},\mathcal{B}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{p\times q}$ are the constraint matrices and $\mathbf{\dot{B}}\in\mathbb{H}^{p\times q}$. Specifically, if $f({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma),g(\mathbf{\dot{W}})=p(\mathbf{\dot{W}})$, $\mathcal{B}=-\mathcal{I}$, $\mathcal{C}=\mathcal{W}$, and $\mathbf{\dot{B}}=0$, then Problem (21) reduces to the SLRQA in (2), where $\mathcal{I}$ denotes the identity operator. The convergence of nonconvex ADMM and its variants has been extensively studied in literatures and the global convergence of PL-ADMM is established.
However, SLRQA-NF does not satisfy a widely-used assumption in existing ADMM-type algorithms. Specifically, SLRQA-NF can be formulated as
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}}$ $f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}}),$ (11)
$\mathrm{s.t.}$ $\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}})=\mathbf{\dot{B}}_{1},$
$\qquad\quad~{}\mathcal{B}_{2}(\mathbf{\dot{W}})=\mathbf{\dot{B}}_{2},$
where $f$ and $g$ are both continuously differentiable, $\mathcal{B}_{1},\mathcal{C}_{1}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{p_{1}\times q_{1}}$, $\mathcal{B}_{2}:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{p_{2}\times q_{2}}$ are given linear maps, $\mathbf{\dot{B}}_{1}\in\mathbb{H}^{p_{1}\times q_{1}}$, and $\mathbf{\dot{B}}_{2}\in\mathbb{H}^{p_{2}\times q_{2}}$. Specifically, if $f({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma),$ $g(\mathbf{\dot{W}})=p(\mathbf{\dot{W}})$, $\mathcal{B}_{1}=-\mathcal{I}$, $\mathcal{C}_{1}=\mathcal{I}$, $\mathcal{B}_{2}=\mathcal{P}_{\Omega}\mathcal{A}\mathcal{W}^{\#}$, $\mathbf{\dot{B}}_{1}=0$, and $\mathbf{\dot{B}}_{2}=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})$, then Problem (11) reduces to the SLRQA-NF in (4). Note that the linear constraints in Problem (11) can be viewed as those in (21) if the two constraints are considered as one, that is $\mathcal{C}=\begin{pmatrix}\mathcal{C}_{1}\\ 0\end{pmatrix}$, $\mathcal{B}=\begin{pmatrix}\mathcal{B}_{1}\\ \mathcal{B}_{2}\end{pmatrix}$, and $\mathbf{\dot{B}}=\begin{pmatrix}\mathbf{\dot{B}}_{1}\\ \mathbf{\dot{B}}_{2}\end{pmatrix}$. In the convergence analysis of existing ADMM-type algorithms for problems with constraints as in (21), the assumption
[…]
which we refer to as “range assumption”, are required [35, 43, 27, 19, 24]. The authors in [35] show the tightness of the “range assumption” in the sense that there exists a problem such that the “range assumption” does not hold while the ADMM algorithm for this problem diverges. This tightness does not prevent a weaker version of the “range assumption” that guarantees convergence of ADMM. Since the constraints in (11) do not satisfy the “range assumption”, we propose an alternative weaker assumption that guarantees global convergence of an ADMM-type algorithm.